# Discrete-state Continuous-time Diffusion for Graph Generation

### v1

@@ 1 Introduction
Some earlier attempts at graph diffusion models treat the graph data in a continuous state space by viewing the graph topology and features as continuous variables [48]. Such a formulation departs from the discrete nature of graph-structured data; e.g., topological sparsity is lost and the discretization in the generation process requires extra hyper-parameters. DiGress [62] is the first work applying discrete-state diffusion models to graph generation tasks and is the current state-of-the-art graph diffusion generative model. However, DiGress is defined in the discrete time space whose generation is inflexible. This is because, its number of sampling steps must match the number of forward diffusion steps, which is a fixed hyperparameter after the model finishes training. A unique advantage of the continuous-time diffusion models [59, 27] lies in their flexible sampling process, and its simulation complexity is proportional to the number of sampling steps, which is determined by the step size of various numerical approaches (e.g., $\tau$-leaping [14, 7, 60]) and decoupled from the training of models. Thus, a discrete-state continuous-time diffusion model is highly desirable for graph generation tasks.
@@
DisCo bears several desirable properties and advantages. First, despite its simplicity, the training objective has a rigorously proved connection to the sampling error. Second, its formulation includes a parametric graph-to-graph mapping, named backbone model, whose input-output architecture is shared between DisCo and DiGress. Therefore, the graph transformer (GT)-based backbone model [46] from DiGress can be seamlessly plugged into DisCo. Third, a concise message-passing neural network backbone model is explored with DisCo, which is much simpler than the GT backbone and performs competitively empirically. Last but not least, our analyses show that the forward and reverse diffusion process in DisCo can retain the permutation-equivariant/invariant properties for both its training loss and sampling distribution; they are critical and practical inductive biases on graph data.
Our comprehensive experiments on plain and molecule graphs show that DisCo can obtain competitive or superior performance against state-of-the-art graph generative models and provide additional sampling flexibility. Our main contributions are summarized as follows.
  • Model. We propose the first discrete-state continuous-time graph diffusion model, DisCo. Also, we utilize the successful graph-to-graph neural network architecture from DiGress and further explore a new lightweight backbone model with decent efficacy.
@@
  • Experiment. Extensive experiments evaluate the empirical performance of DisCo.
@@ 2.1 Discrete-State Continuous-time Diffusion Models
A $D$-dimensional discrete state space is represented as $\mathcal{X}=\{1,\dots,C\}^{D}$. A continuous-time Markov Chain (CTMC) $\{\mathbf{x}_{t}=[x_{t}^{1},\dotsm x_{t}^{D}]\}_{t\in[0,T]}$ is characterized by its (time-dependent) rate matrix $\mathbf{R}_{t}\in\mathbb{R}^{|\mathcal{X}|\times|\mathcal{X}|}$. Here $\mathbf{x}_{t}$ is the state at the time step $t$. The transition probability $q_{t|s}$ between the time interval $s$ to $t$ satisfies the Kolmogorov forward equation, for $s<t$,
@@
The marginal distribution can be represented as
$q_{t}(\mathbf{x}_{t})=\sum_{\mathbf{x}_{0}\in\mathcal{X}}q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})\pi_{\texttt{data}}(\mathbf{x}_{0}),$ (2)
where $\pi_{\texttt{data}}(\mathbf{x}_{0})$ is the data distribution. If the CTMC is defined in time interval $[0,T]$ and if the rate matrix $\mathbf{R}_{t}$ is well-designed, the final distribution $q_{T}(\mathbf{x}_{T})$ can be close to a tractable reference distribution $\pi_{\texttt{ref}}(\mathbf{x}_{T})$, e.g., uniform distribution. We notate the reverse stochastic process as $\tilde{\mathbf{x}}_{t}=\mathbf{x}_{T-t}$, A well-known fact (e.g., Section 5.9 in  [53]) is that the reverse process $\{\tilde{\mathbf{x}}_{t}\}_{t\in[0,T]}$ is also a CTMC. It is characterized by the reverse rate matrix defined as
$\tilde{\mathbf{R}}_{t}(\mathbf{x},\mathbf{y})=\frac{q(\mathbf{y})}{q(\mathbf{x})}\mathbf{R}_{t}(\mathbf{y},\mathbf{x}).$ (3)
The goal of the CTMC-based diffusion models is an accurate estimation of the reverse rate matrix $\tilde{\mathbf{R}}_{t}$ so that new data can be generated by sampling the reference distribution $\pi_{\texttt{ref}}$ and then simulating the reverse CTMC [12, 13, 14, 1]. However, the complexity of the rate matrix is prohibitively high because there are $C^{D}$ possible states. A reasonable simplification is to factorize the process over dimensions [7, 60, 62, 2]. Specifically, the forward process is factorized as $q_{t|s}(\mathbf{x}_{t}|\mathbf{x}_{s})=\prod_{d=1}^{D}q_{t|s}(x_{t}^{d}|x_{s}^{d})$, for $s<t$. Then, the forward diffusion of each dimension is independent and is governed by dimension-specific forward rate matrices $\{\mathbf{R}_{t}^{d}\}_{d=1}^{D}$. With such a factorization, the goal is to estimate the dimension-specific reverse rate matrices $\{\tilde{\mathbf{R}}^{d}_{t}\}_{d=1}^{D}$.
The reverse rate of one dimension is represented as $\tilde{\mathbf{R}}_{t}^{d}(x^{d},y^{d})=\sum_{x_{0}^{d}}\mathbf{R}^{d}_{t}(y^{d},x^{d})\frac{q_{t|0}(y^{d}|x^{d}_{0})}{q_{t|0}(x^{d}|x^{d}_{0})}q_{0|t}(x^{d}_{0}|\mathbf{x})$. Campbell et al. [7] estimate $q_{0|t}(x^{d}_{0}|\mathbf{x})$ via a neural network $p_{\theta}$ such that $p_{\theta}(x^{d}_{0}|\mathbf{x},t)\approx q_{0|t}(x^{d}_{0}|\mathbf{x})$; Sun et al. [60] propose another singleton conditional distribution-based objective $\frac{p_{\theta}(y^{d}|\mathbf{x}^{\setminus d},t)}{p_{\theta}(x^{d}|\mathbf{x}^{\setminus d},t)}\approx\frac{q(y^{d}|\mathbf{x}^{\setminus d})}{q(x^{d}|\mathbf{x}^{\setminus d})}$ whose rationale is Brook’s Lemma [4, 41].
@@ 2.2 Graph Generation and Notations
We study the generation of graphs with categorical node and edge attributes. A graph with $n$ nodes is represented by its edge type matrix and node type vector: $\mathcal{G}=(\mathbf{E},\mathbf{F})$, where $\mathbf{E}=\{e^{(i,j)}\}_{i,j\in\mathbb{N}^{+}_{\leq n}}\in\{1,\dots,a,a+1\}^{n\times n}$, $\mathbf{F}=\{f^{i}\}_{i\in\mathbb{N}^{+}_{\leq n}}\in\{1,\dots,b\}^{n}$, $a$ is the number of edge types, and $b$ is the number of node types. Notably, the absence of an edge is viewed as a special edge type, so that in total there are $(a+1)$ edge types. The problem we study is defined as follows.
Problem 2.1.
Graph Generation
Given: $N$ graphs $\{\mathcal{G}^{i}\}_{i\in\mathbb{N}^{+}_{\leq N}}$ from an inaccessible graph data distribution $\mathfrak{G}$.
Find: $M$ generated graphs $\{\mathcal{G}^{i}\}_{i\in\mathbb{N}^{+}_{\leq M}}$ from the distribution $\mathfrak{G}$.
@@ 3 Method
Figure 2: An overview of discrete-state continuous-time graph diffusion. A transition can happen at any time in $[0,T]$.
This section presents the proposed discrete-state continuous-time graph diffusion model, DisCo. An overview of DisCo is presented in Figure 2. Section 3.1 introduces the necessity to factorize the diffusion process and Section 3.2 details the forward process. Our training objective and its connection to sampling are introduced in Sections 3.3 and 3.4, respectively. Last but not least, a specific neural architecture of the graph-to-graph backbone model and its properties regarding the permutation of node ordering are introduced in Sections 3.5 and 3.6, respectively.
All proofs are in Appendix.
@@ 3.1 Factorized Discrete Graph Diffusion Process
$\tilde{\mathbf{R}}_{t}(\mathcal{G},\bar{\mathcal{G}})$ $=\sum_{i}A_{t}^{i}\sum_{f_{0}^{i}}\frac{q_{t|0}(\bar{f}^{i}|f^{i}_{0})}{q_{t|0}(f^{i}|f^{i}_{0})}q_{0|t}(f^{i}_{0}|\mathcal{G})+\sum_{i,j}B_{t}^{(i,j)}\sum_{e_{0}^{(i,j)}}\frac{q_{t|0}(\bar{e}^{(i,j)}|e^{(i,j)}_{0})}{q_{t|0}(e^{(i,j)}|e^{(i,j)}_{0})}q_{0|t}(e^{(i,j)}_{0}|\mathcal{G}),$ (6)
@@
Note that this factorization itself is not our contribution but a necessary part of our framework, so we mention it here for completeness. Its full derivation is in Appendix - Section A. Next, we detail the design of forward rate matrices and the estimation of reverse rate matrices.
@@ 3.2 Forward Process
The forward process for nodes and edges converges to uniform distributions if $\mathbf{R}_{f}=\mathbf{1}\mathbf{1}^{\top}-b\mathbf{I}$ and $\mathbf{R}_{e}=\mathbf{1}\mathbf{1}^{\top}-(a+1)\mathbf{I}$. The forward process for nodes and edges converges to marginal distributions $\mathbf{m}_{f}$ and $\mathbf{m}_{e}$ if $\mathbf{R}_{f}=\mathbf{1}\mathbf{m}_{f}^{\top}-\mathbf{I}$ and $\mathbf{R}_{e}=\mathbf{1}\mathbf{m}_{e}^{\top}-\mathbf{I}$. For both cases, $\mathbf{R}_{e}\in\mathbb{R}^{(a+1)\times(a+1)}$, $\mathbf{R}_{f}\in\mathbb{R}^{b\times b}$, $\mathbf{1}$ is an all-one column vector, and $\mathbf{I}$ is an identity matrix.
@@
1:Given a training graph $\mathcal{G}_{0}=(\{f_{0}^{i}\},\{e_{0}^{(i,j)}\})$
2:Sample $t\sim\mathcal{U}_{(0,T)}$
@@
4:Predict the clean graph $\hat{\mathcal{G}}_{0}=(\{\hat{f}_{0}^{i}\},\{\hat{e}_{0}^{(i,j)}\})\leftarrow\Big{(}\{p^{\theta}_{0|t}(f^{i}|\mathcal{G}_{t})\},\{p^{\theta}_{0|t}(e^{(i,j)}|\mathcal{G}_{t})\}\Big{)}$, given $\mathcal{G}_{t}$
5:Compute cross-entropy loss between $\mathcal{G}_{0}$ and $\hat{\mathcal{G}}_{0}$ based on Eq. (9) and update $\theta$
@@ 3.3 Parameterization and Optimization Objective
Next, we introduce the estimation and parameterization of the reverse process from its motivation. The reverse process is essentially determined by the reverse rate matrix $\tilde{\mathbf{R}}_{t}$ in Eq. (6), whose computation needs $q_{0|t}(f^{i}_{0}|\mathcal{G})$ and $q_{0|t}(e^{(i,j)}_{0}|\mathcal{G})$; they are expensive to estimate exactly because according to Bayes’ rule, $p_{t}(\mathcal{G})$ is needed, whose computation needs to enumerate all the given graphs: $p_{t}(\mathcal{G})=\sum_{\mathcal{G}_{0}}q_{t|0}(\mathcal{G}|\mathcal{G}_{0})\pi_{\texttt{data}}(\mathcal{G}_{0})$.
Thus, we propose parameterizing the reverse transition probabilities via a neural network $\theta$ whose specific architecture is introduced in Section 3.5. The terms $\{q_{0|t}(f^{i}_{0}|\mathcal{G})\}_{i\in\mathbb{N}^{+}_{\leq n}}$ and $\{q_{0|t}(e^{(i,j)}_{0}|\mathcal{G})\}_{i,j\in\mathbb{N}^{+}_{\leq n}}$ in Eq. (6) are replaced with the parameterized $\{p^{\theta}_{0|t}(f^{i}|\mathcal{G})\}_{i\in\mathbb{N}^{+}_{\leq n}}$ and $\{p^{\theta}_{0|t}(e^{(i,j)}|\mathcal{G})\}_{i,j\in\mathbb{N}^{+}_{\leq n}}$. Thus, a parameterized reverse rate matrix $\tilde{\mathbf{R}}_{\theta,t}(\mathcal{G},\bar{\mathcal{G}})$ is represented as
where $\tilde{\mathbf{R}}^{i}_{\theta,t}(f^{i},\bar{f}^{i})=A_{t}^{i}\sum_{f_{0}^{i}}\frac{q_{t|0}(\bar{f}^{i}|f^{i}_{0})}{q_{t|0}(f^{i}|f^{i}_{0})}p^{\theta}_{0|t}(f^{i}_{0}|\mathcal{G})$, $\tilde{\mathbf{R}}^{(i,j)}_{\theta,t}(e^{(i,j)},\bar{e}^{(i,j)})=B_{t}^{(i,j)}\sum_{e_{0}^{(i,j)}}\frac{q_{t|0}(\bar{e}^{(i,j)}|e^{(i,j)}_{0})}{q_{t|0}(e^{(i,j)}|e^{(i,j)}_{0})}p^{\theta}_{0|t}(e^{(i,j)}_{0}|\mathcal{G})$, and the remaining notations are the same as Eq. (6). Note that all the above terms $\{p^{\theta}_{0|t}(f^{i}|\mathcal{G})\}_{i\in\mathbb{N}^{+}_{\leq n}}$ and $\{p^{\theta}_{0|t}(e^{(i,j)}|\mathcal{G})\}_{i,j\in\mathbb{N}^{+}_{\leq n}}$ can be viewed together as a graph-to-graph mapping $\theta:\mathfrak{G}\to\mathfrak{G}$, whose input is the noisy graph $\mathcal{G}_{t}$ and its output is the predicted clean graph probabilities, concretely, the node/edge type probabilities of all the nodes and edges.
Intuitively, the discrepancy between the groundtruth reverse rate matrix $\tilde{\mathbf{R}}_{t}$ (from Eq. (6)) and the parametric reverse rate matrix $\tilde{\mathbf{R}}_{\theta,t}$ should be small. Our following analysis (Theorem 3.3) establishes an upper bound of such a discrepancy, where the node and edge estimated probability vector (sum is $1$) is notated as $\hat{f}^{i}_{0}=[p^{\theta}_{0|t}(f^{i}=1|\mathcal{G}_{t}),\dots,p^{\theta}_{0|t}(f^{i}=b|\mathcal{G}_{t})]^{\top}\in[0,1]^{b}$ and $\hat{e}^{(i,j)}_{0}=[p^{\theta}_{0|t}(e^{(i,j)}=1|\mathcal{G}_{t}),\dots,p^{\theta}_{0|t}(e^{(i,j)}=a+1|\mathcal{G}_{t})]^{\top}\in[0,1]^{a+1}$.
@@
The bound in Theorem 3.3 is tight, i.e., the right-hand side of Eq. (8) is 00, whenever $\hat{f}_{0}^{i}=q_{0|t}(f_{0}^{i}|\mathcal{G}_{t}),\forall i$ and $\hat{e}_{0}^{(i,j)}=q_{0|t}(e_{0}^{(i,j)}|\mathcal{G}_{t}),\forall i,j$. Guided by Theorem 3.3, we (1) take expectation over all the time $t$ by sampling $t$ from a uniform distribution and (2) simplify the right-hand side of Eq. (8) and use the unweighted cross-entropy (CE) loss as our training objective:
A step-by-step training algorithm is presented in Algorithm 1. It is worth mentioning that the above CE loss has been used in some diffusion models (e.g., [2, 7]) but lacks a good motivation, especially in the continuous-time setting. Our analysis motivates it based on the rate matrix discrepancy, as a unique contribution of this paper.

### v2

@@ 1 Introduction
Some earlier attempts at graph diffusion models treat the graph data in a continuous state space by viewing the graph topology and features as continuous variables [56]. Such a formulation departs from the discrete nature of graph-structured data; e.g., topological sparsity is lost and the discretization in the generation process requires extra hyper-parameters. DiGress [73] is one of the early efforts applying discrete-state diffusion models to graph generation tasks and is the current state-of-the-art graph diffusion generative model. However, DiGress is defined in the discrete time space whose generation is inflexible. This is because, its number of sampling steps must match the number of forward diffusion steps, which is a fixed hyperparameter after the model finishes training. A unique advantage of the continuous-time diffusion models [70, 32] lies in their flexible sampling process, and its simulation complexity is proportional to the number of sampling steps, determined by the step size of various numerical approaches (e.g., $\tau$-leaping [18, 8, 71]) and decoupled from the models’ training. Thus, a discrete-state continuous-time diffusion model is highly desirable for graph generation tasks.
@@
DisCo bears several desirable properties and advantages. First, despite its simplicity, the training objective has a rigorously proved connection to the sampling error. Second, its formulation includes a parametric graph-to-graph mapping, named backbone model, whose input-output architecture is shared between DisCo and DiGress. Therefore, the graph transformer (GT)-based backbone model [54] from DiGress can be seamlessly plugged into DisCo. Third, a concise message-passing neural network backbone model is explored with DisCo, which is simpler than the GT backbone and has decent empirical performance. Last but not least, our analyses show that the forward and reverse diffusion process in DisCo can retain the permutation-equivariant/invariant properties for its training loss and sampling distribution, both of which are critical and practical inductive biases on graph data.
Comprehensive experiments on plain and molecule graphs show that DisCo can obtain competitive or superior performance against state-of-the-art graph generative models and provide additional sampling flexibility. Our main contributions are summarized:
  • Model. We propose the first discrete-state continuous-time graph diffusion model, DisCo. We utilize the successful graph-to-graph neural network architecture from DiGress and further explore a new lightweight backbone model with decent efficacy.
@@
  • Experiment. Extensive experiments validate the empirical performance of DisCo.
@@ 2.1 Discrete-State Continuous-time Diffusion Models
A $D$-dimensional discrete state space is represented as $\mathcal{X}=\{1,\dots,C\}^{D}$. A continuous-time Markov Chain (CTMC) $\{\mathbf{x}_{t}=[x_{t}^{1},\dotsm x_{t}^{D}]\}_{t\in[0,T]}$ is characterized by its (time-dependent) rate matrix $\mathbf{R}_{t}\in\mathbb{R}^{|\mathcal{X}|\times|\mathcal{X}|}$. Here $\mathbf{x}_{t}$ is the state at the time step $t$. The transition probability $q_{t|s}$ between from time $s$ to $t$ satisfies the Kolmogorov forward equation, for $s<t$,
@@
The marginal distribution can be represented as $q_{t}(\mathbf{x}_{t})=\sum_{\mathbf{x}_{0}\in\mathcal{X}}q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})\pi_{\texttt{data}}(\mathbf{x}_{0})$ where $\pi_{\texttt{data}}(\mathbf{x}_{0})$ is the data distribution. If the CTMC is defined in time interval $[0,T]$ and if the rate matrix $\mathbf{R}_{t}$ is well-designed, the final distribution $q_{T}(\mathbf{x}_{T})$ can be close to a tractable reference distribution $\pi_{\texttt{ref}}(\mathbf{x}_{T})$, e.g., uniform distribution. We notate the reverse stochastic process as $\tilde{\mathbf{x}}_{t}=\mathbf{x}_{T-t}$; a well-known fact (e.g., Section 5.9 in  [63]) is that the reverse process $\{\tilde{\mathbf{x}}_{t}\}_{t\in[0,T]}$ is also a CTMC, characterized by the reverse rate matrix: $\tilde{\mathbf{R}}_{t}(\mathbf{x},\mathbf{y})=\frac{q(\mathbf{y})}{q(\mathbf{x})}\mathbf{R}_{t}(\mathbf{y},\mathbf{x})$. The goal of the CTMC-based diffusion models is an accurate estimation of the reverse rate matrix $\tilde{\mathbf{R}}_{t}$ so that new data can be generated by sampling the reference distribution $\pi_{\texttt{ref}}$ and then simulating the reverse CTMC [16, 17, 18, 1]. However, the complexity of the rate matrix is prohibitively high because there are $C^{D}$ possible states. A reasonable simplification is to factorize the process over dimensions [8, 71, 73, 2]. Specifically, the forward process is factorized as $q_{t|s}(\mathbf{x}_{t}|\mathbf{x}_{s})=\prod_{d=1}^{D}q_{t|s}(x_{t}^{d}|x_{s}^{d})$, for $s<t$. Then, the forward diffusion of each dimension is independent and is governed by dimension-specific forward rate matrices $\{\mathbf{R}_{t}^{d}\}_{d=1}^{D}$. With such a factorization, the goal is to estimate the dimension-specific reverse rate matrices $\{\tilde{\mathbf{R}}^{d}_{t}\}_{d=1}^{D}$.
The dimension-specific reverse rate is represented as $\tilde{\mathbf{R}}_{t}^{d}(x^{d},y^{d})=\sum_{x_{0}^{d}}\mathbf{R}^{d}_{t}(y^{d},x^{d})\frac{q_{t|0}(y^{d}|x^{d}_{0})}{q_{t|0}(x^{d}|x^{d}_{0})}q_{0|t}(x^{d}_{0}|\mathbf{x})$. Campbell et al. [8] estimate $q_{0|t}(x^{d}_{0}|\mathbf{x})$ via a neural network $p_{\theta}$ such that $p_{\theta}(x^{d}_{0}|\mathbf{x},t)\approx q_{0|t}(x^{d}_{0}|\mathbf{x})$; Sun et al. [71] propose another singleton conditional distribution-based objective $\frac{p_{\theta}(y^{d}|\mathbf{x}^{\setminus d},t)}{p_{\theta}(x^{d}|\mathbf{x}^{\setminus d},t)}\approx\frac{q(y^{d}|\mathbf{x}^{\setminus d})}{q(x^{d}|\mathbf{x}^{\setminus d})}$ whose rationale is Brook’s Lemma [5, 49].
@@ 2.2 Graph Generation and Notations
We study the graphs with categorical node and edge attributes. A graph with $n$ nodes is represented by its edge type matrix and node type vector: $\mathcal{G}=(\mathbf{E},\mathbf{F})$, where $\mathbf{E}=(e^{(i,j)})_{i,j\in\mathbb{N}^{+}_{\leq n}}\in\{1,\dots,a+1\}^{n\times n}$, $\mathbf{F}=(f^{i})_{i\in\mathbb{N}^{+}_{\leq n}}\in\{1,\dots,b\}^{n}$, $a$ and $b$ are the numbers of node and edge types, respectively. Notably, the absence of an edge is viewed as a special edge type, so there are $(a+1)$ edge types in total. The problem we study is graph generation where $N$ graphs $\{\mathcal{G}^{i}\}_{i\in\mathbb{N}^{+}_{\leq N}}$ from an inaccessible graph data distribution $\mathfrak{G}$ are given and we aim to generate $M$ graphs $\{\mathcal{G}^{i}\}_{i\in\mathbb{N}^{+}_{\leq M}}$ from $\mathfrak{G}$.
@@ 3 Method
Figure 2: An overview of DisCo. A transition can happen at any time in $[0,T]$.
This section presents the proposed discrete-state continuous-time graph diffusion model, DisCo whose overview is Figure 2. Section 3.1 introduces the necessity to factorize the diffusion process and Section 3.2 details the forward process. Our training objective and its connection to sampling are introduced in Sections 3.3 and 3.4, respectively. Last but not least, a specific neural architecture of the graph-to-graph backbone model and its properties regarding the permutation of node ordering are introduced in Sections 3.5 and 3.6, respectively. All proofs are in Appendix.
@@ 3.1 Factorized Discrete Graph Diffusion Process
$\tilde{\mathbf{R}}_{t}(\mathcal{G},\bar{\mathcal{G}})$ $=\sum_{i}A_{t}^{i}\sum_{f_{0}^{i}}\frac{q_{t|0}(\bar{f}^{i}|f^{i}_{0})}{q_{t|0}(f^{i}|f^{i}_{0})}q_{0|t}(f^{i}_{0}|\mathcal{G})+\sum_{i,j}B_{t}^{(i,j)}\sum_{e_{0}^{(i,j)}}\frac{q_{t|0}(\bar{e}^{(i,j)}|e^{(i,j)}_{0})}{q_{t|0}(e^{(i,j)}|e^{(i,j)}_{0})}q_{0|t}(e^{(i,j)}_{0}|\mathcal{G})$ (4)
@@
Note that this factorization itself is not our contribution but a necessary part of our framework, so we mention it here for completeness. Its full derivation is in Appendix - Section A. Next, we detail the design of forward rate matrices.
@@ 3.2 Forward Process
The forward processes for nodes and edges converge to uniform distributions if $\mathbf{R}_{f}=\mathbf{1}\mathbf{1}^{\top}-b\mathbf{I}$ and $\mathbf{R}_{e}=\mathbf{1}\mathbf{1}^{\top}-(a+1)\mathbf{I}$; they converge to marginal distributions $\mathbf{m}_{f}$ and $\mathbf{m}_{e}$ if $\mathbf{R}_{f}=\mathbf{1}\mathbf{m}_{f}^{\top}-\mathbf{I}$ and $\mathbf{R}_{e}=\mathbf{1}\mathbf{m}_{e}^{\top}-\mathbf{I}$. $\mathbf{1}$ is an all-one vector and $\mathbf{I}$ is an identity matrix.
@@
1:A training graph $\mathcal{G}_{0}=(\{f_{0}^{i}\},\{e_{0}^{(i,j)}\})$ is given.
@@
3:Predict the clean graph $\hat{\mathcal{G}}_{0}=(\{\hat{f}_{0}^{i}\},\{\hat{e}_{0}^{(i,j)}\})\leftarrow\Big{(}\{p^{\theta}_{0|t}(f^{i}|\mathcal{G}_{t})\},\{p^{\theta}_{0|t}(e^{(i,j)}|\mathcal{G}_{t})\}\Big{)}$, given $\mathcal{G}_{t}$.
4:Compute cross-entropy loss between $\mathcal{G}_{0}$ and $\hat{\mathcal{G}}_{0}$ based on Eq. (6) and update $\theta$.
@@ 3.3 Parameterization and Optimization Objective
Next, we introduce the estimation of the reverse process from its motivation. The reverse process is essentially determined by the reverse rate matrix $\tilde{\mathbf{R}}_{t}$ in Eq. (4), whose computation needs $q_{0|t}(f^{i}_{0}|\mathcal{G})$ and $q_{0|t}(e^{(i,j)}_{0}|\mathcal{G})$, $\forall i,j$; their exact estimation is expensive because according to Bayes’ rule, $p_{t}(\mathcal{G})$ is needed, whose computation needs to enumerate all the given graphs: $p_{t}(\mathcal{G})=\sum_{\mathcal{G}_{0}}q_{t|0}(\mathcal{G}|\mathcal{G}_{0})\pi_{\texttt{data}}(\mathcal{G}_{0})$.
Thus, we propose parameterizing the reverse transition probabilities via a neural network $\theta$ whose specific architecture is introduced in Section 3.5. The terms $\{q_{0|t}(f^{i}_{0}|\mathcal{G})\}_{i\in\mathbb{N}^{+}_{\leq n}}$ and $\{q_{0|t}(e^{(i,j)}_{0}|\mathcal{G})\}_{i,j\in\mathbb{N}^{+}_{\leq n}}$ in Eq. (4) are replaced with the parameterized $\{p^{\theta}_{0|t}(f^{i}|\mathcal{G})\}_{i\in\mathbb{N}^{+}_{\leq n}}$ and $\{p^{\theta}_{0|t}(e^{(i,j)}|\mathcal{G})\}_{i,j\in\mathbb{N}^{+}_{\leq n}}$. Thus, a parameterized reverse rate matrix $\tilde{\mathbf{R}}_{\theta,t}(\mathcal{G},\bar{\mathcal{G}})$ is represented as $\tilde{\mathbf{R}}_{\theta,t}(\mathcal{G},\bar{\mathcal{G}})=\sum_{i}\tilde{\mathbf{R}}^{i}_{\theta,t}(f^{i},\bar{f}^{i})+\sum_{i,j}\tilde{\mathbf{R}}^{(i,j)}_{\theta,t}(e^{(i,j)},\bar{e}^{(i,j)})$ where $\tilde{\mathbf{R}}^{i}_{\theta,t}(f^{i},\bar{f}^{i})=A_{t}^{i}\sum_{f_{0}^{i}}\frac{q_{t|0}(\bar{f}^{i}|f^{i}_{0})}{q_{t|0}(f^{i}|f^{i}_{0})}p^{\theta}_{0|t}(f^{i}_{0}|\mathcal{G})$, $\tilde{\mathbf{R}}^{(i,j)}_{\theta,t}(e^{(i,j)},\bar{e}^{(i,j)})=B_{t}^{(i,j)}\sum_{e_{0}^{(i,j)}}\frac{q_{t|0}(\bar{e}^{(i,j)}|e^{(i,j)}_{0})}{q_{t|0}(e^{(i,j)}|e^{(i,j)}_{0})}p^{\theta}_{0|t}(e^{(i,j)}_{0}|\mathcal{G})$, and the remaining notations are the same as Eq. (4). Note that all the terms $\{p^{\theta}_{0|t}(f^{i}|\mathcal{G})\}_{i\in\mathbb{N}^{+}_{\leq n}}$ and $\{p^{\theta}_{0|t}(e^{(i,j)}|\mathcal{G})\}_{i,j\in\mathbb{N}^{+}_{\leq n}}$ can be viewed together as a graph-to-graph mapping $\theta:\mathfrak{G}\mapsto\mathfrak{G}$, whose input is the noisy graph $\mathcal{G}_{t}$ and its output is the predicted clean graph probabilities, concretely, the node/edge type probabilities of all the nodes and edges.
Intuitively, the discrepancy between the groundtruth $\tilde{\mathbf{R}}_{t}$ (from Eq. (4)) and the parametric $\tilde{\mathbf{R}}_{\theta,t}$ should be small. Theorem 3.3 establishes a cross-entropy (CE)-based upper bound of such a discrepancy, where the estimated probability vectors (sum is $1$) are notated as $\hat{f}^{i}_{0}=[p^{\theta}_{0|t}(f^{i}=1|\mathcal{G}_{t}),\dots,p^{\theta}_{0|t}(f^{i}=b|\mathcal{G}_{t})]^{\top}\in[0,1]^{b}$ and $\hat{e}^{(i,j)}_{0}=[p^{\theta}_{0|t}(e^{(i,j)}=1|\mathcal{G}_{t}),\dots,p^{\theta}_{0|t}(e^{(i,j)}=a+1|\mathcal{G}_{t})]^{\top}\in[0,1]^{a+1}$.
@@
The bound in Theorem 3.3 is tight, i.e., the right-hand side of Eq. (5) is 00, whenever $\hat{f}_{0}^{i}=q_{0|t}(f_{0}^{i}|\mathcal{G}_{t}),\forall i$ and $\hat{e}_{0}^{(i,j)}=q_{0|t}(e_{0}^{(i,j)}|\mathcal{G}_{t}),\forall i,j$. Guided by Theorem 3.3, we (1) take expectation of $t$ by sampling $t$ from a uniform distribution $t\sim\mathcal{U}_{(0,T)}$ and (2) simplify the right-hand side of Eq. (5) by using the unweighted CE loss as our training objective:
$\min_{\theta}\ T\mathbb{E}_{t}\mathbb{E}_{\mathcal{G}_{0}}\mathbb{E}_{q_{t|0}(\mathcal{G}_{t}|\mathcal{G}_{0})}\Big{[}\sum_{i}\mathcal{L}_{\mathtt{CE}}(\mathtt{One\text{-}Hot}(f_{0}^{i}),\hat{f}_{0}^{i})+\sum_{i,j}\mathcal{L}_{\mathtt{CE}}(\mathtt{One\text{-}Hot}(e_{0}^{(i,j)}),\hat{e}_{0}^{(i,j)})\Big{]}$ (6)
A step-by-step training algorithm is in Algorithm 1. Note that the above CE loss has been used in some diffusion models (e.g., [2, 8]) but lacks a good motivation, especially in the continuous-time setting. We motivate it based on the rate matrix discrepancy, as a unique contribution of this paper.